OpenStudy (anonymous):

A lottery game has balls number 1 throgh 25. What is the probability of selecting an even numbered ball or a 19?

OpenStudy (campbell_st):

P(Even) = 12/25 P(19) = 1/25 P(even or 19) = 12/25 + 1/25

OpenStudy (anonymous):

i think cam.. is right